a gas occupies 0.6 litres under a pressure of 0.92 bar. under what pressure will the volume of gas be reduced by 20% of its original volume ? ( temp is constant)

Grade:11

4 Answers

Naveen Gunturu
14 years ago

is the answer 4.6?

anurag singhal
14 years ago

at a constant temperature   P is infersely  proportionsl  to   1/V

PV=comstant

at same temperature apply P1V1 = P2V2

thus u will get P as 1.15 bar

trishit thakur
14 years ago

P1V1=P2V2
.92*.6= .12*P2

P2= 4.6 bar :answer

askiitian.expert- chandra sekhar
14 years ago

Hi Tejas,

V1=0.6 lit

V2=V1-V1*20/100

    = V1(1-20/100)=V1(80/100)

P1=0.92 bar

Boyle's law: P1*V1=P2*V2

                 0.92*V1=P2*V1*0.8

                  P2=0.92/0.8=1.15 bar